# Zero-Shot Transfer Capabilities of the Sundial Foundation Model for Leaf Area Index Forecasting

### v1

@@ Abstract
This work investigates the zero-shot forecasting capability of time-series foundation models for Leaf Area Index (LAI) forecasting in agricultural monitoring. Using the HiQ dataset (U.S., 2000-2022), we systematically compare statistical baselines, a fully supervised LSTM, and the Sundial foundation model under multiple evaluation protocols. We find that Sundial, in the zero-shot setting, can outperform a fully trained LSTM provided that the input context window is sufficiently long—specifically, when covering more than one or two full seasonal cycles. This demonstrates, for the first time, that a general-purpose foundation model can surpass specialized supervised models on remote-sensing time series prediction without any task-specific tuning. These results highlight the strong potential of pretrained time-series foundation models to serve as effective plug-and-play forecasters in agricultural and environmental applications.
@@ Introduction
Leaf Area Index (LAI) is a critical biophysical variable widely used in agricultural and ecological monitoring, quantifying canopy structure and influencing crop growth. Beyond its central role in agronomic decision-making and yield estimation, accurate LAI forecasting is vital for understanding land-surface feedbacks, as vegetation significantly influences regional climate through biogeophysical processes (e.g., albedo and roughness) and biogeochemical cycles (alessandri2017multi). However, practical forecasting is challenged by the diversity of cropping systems, varying environmental conditions, and the limited availability of labeled data, motivating a need for robust and transferable models capable of generalizing across locations and years. Furthermore, traditional supervised models trained on heterogeneous datasets tend to learn ”global average” patterns, often struggling to capture site-specific phenological anomalies without extensive local fine-tuning.
Recent advances in time series foundation models, such as Sundial (liu2025sundial), offer a promising avenue for improving LAI forecasting. Unlike task-specific models, these large, pretrained networks leverage in-context learning, treating historical time series as ”prompts” to deduce underlying dynamics. This paradigm has the potential to adapt to new tasks and domains more efficiently than traditional supervised approaches.
@@
  • We analyze the context window threshold where zero-shot performance overtakes the supervised baseline, highlighting the distinct mechanisms at play: instance-level adaptation in foundation models versus dataset-level generalization in LSTMs.
@@
LAI Time-Series Forecasting
Leaf Area Index (LAI) is a key indicator for monitoring vegetation health, phenology, and long-term ecosystem dynamics. Most LAI forecasting approaches treat the signal as a univariate time series characterized by strong seasonality, low-frequency variation, and limited annotated samples (jiang2010modeling). Classical methods include harmonic models, autoregressive formulations, and exponential-smoothing variants, which capture seasonal cycles but struggle with nonlinear responses to climate and land-cover variability (winters1960forecasting). Recent studies explore data-driven forecasting of vegetation indices, yet robust LAI prediction remains challenging due to heterogeneous temporal patterns and spatial variability across regions (liu2022estimation).
@@
Classical LAI forecasting has long relied on statistical time-series models such as autoregressive (AR) and exponential-smoothing formulations, which are simple, interpretable, and effective for short-term dynamics. However, these models are limited in their ability to capture nonlinear seasonal patterns and interannual variability.
Supervised machine-learning baselines—including random forests (breiman2001random) and gradient boosting (friedman2001greedy; ke2017lightgbm)—can incorporate richer auxiliary features when available, but offer only marginal improvements when restricted to univariate LAI histories. Recurrent neural networks, particularly LSTMs, remain strong supervised baselines for vegetation-index forecasting and are widely used in agronomic and ecological applications (hochreiter1997long; liu2022machine). Despite their capacity to model nonlinear temporal dependencies, these methods require substantial labeled data and typically exhibit limited gains when training data is scarce or spatially heterogeneous.
Importantly, none of these approaches leverage cross-domain temporal knowledge. As a result, their generalization under long sequences, sparse supervision, or unseen spatial regions remains constrained—highlighting the value of pretrained time-series foundation models, which can transfer temporal priors learned from large and diverse corpora.
Time-Series Foundation Models
Large pretrained time-series models have recently demonstrated strong zero-shot and few-shot transfer across diverse domains. Models such as TimesFM, and Sundial learn universal temporal representations from mixed-modality corpora spanning climate, sensor networks, and economic indicators (das2024decoder; liu2025sundial; ansari2024chronos). These foundation models can leverage broad temporal priors—seasonality, trends, and multi-scale structure—reducing dependence on in-domain training data and enabling robust downstream forecasting.
@@
We conduct our experiments using the HiQ dataset (yan2023hiq), a large-scale benchmark for Leaf Area Index (LAI) prediction. HiQ reprocessed satellite MODIS LAI product by the spatiotemporal information compositing algorithm to ensure better spatiotemporal consistency. It offers high-quality global LAI time series spanning 2000–2022, yielding broad geographic coverage across diverse crop and ecotype regions. In this study, we use the HiQ 5-km spatial resolution product as the benchmark. Each pixel contains a temporal LAI sequence at an 8-day cadence, resulting in approximately 1050 time steps per location.
@@ Methods
  • Trend: Extrapolates a fitted linear trend from the previous $k$ time points to predict future LAI values (jensen2008discrete).
  • ARIMA: For each test window, we fit an ARIMA(2,1,0)(2,1,0) model (box1970distribution; seabold2010statsmodels) to the observed LAI values to forecast the next value. ARIMA is chosen as a simple yet reasonable baseline for univariate time series; model fitting uses state-space estimation.
  • LSTM: A supervised Long Short-Term Memory model (hochreiter1997long) trained on the entire training set to capture nonlinear dependencies in LAI time series. Our LSTM architecture consists of two stacked LSTM layers with 64 hidden units each, followed by a fully connected output layer. The model is trained using the Adam (kingma2014adam) optimizer with a learning rate of 0.001, batch size of 256, and early stopping based on validation loss.
@@
We formulate the LAI forecasting task as a conditional generation problem. Instead of fine-tuning model weights, we leverage the pre-trained capabilities of Sundial directly (liu2025sundial). The historical input window $X\in\mathbb{R}^{T_{\text{in}}}$ serves as a continuous context prompt, analogous to a text prompt in Large Language Models (LLMs). The model processes this prompt to deduce the underlying phenological dynamics (e.g., seasonality and trend) and generates the future sequence. This approach relies entirely on the model’s in-context learning ability acquired during pre-training, enabling a strictly zero-shot evaluation. We use the publicly available Sundial model checkpoint without any modifications or fine-tuning. For inference, we generate forecasts autoregressively for multi-step horizons.
@@
Metrics are aggregated over 100 randomly sampled locations to ensure computational efficiency. With approximately 100 rolling test windows per pixel, each reported metric represents an average across $\sim$ 10,000 prediction-ground truth pairs.
@@ Discussion
In this work, we conducted the first rigorous evaluation of the zero-shot capabilities of the Sundial time-series foundation model for remote-sensing LAI forecasting. We found that, when provided with a sufficiently long historical context window, Sundial’s zero-shot performance can surpass that of a supervised LSTM model specifically trained for the task, challenging the conventional view that task-specific training guarantees superior performance. Our findings suggest that these general-purpose models already encode powerful temporal patterns and vegetation phenology, enabling them to serve as highly efficient and powerful plug-and-play forecasting tools without the need for any task-specific training.

### v2

@@ Abstract
This work investigates the zero-shot forecasting capability of time series foundation models for Leaf Area Index (LAI) forecasting in agricultural monitoring. Using the HiQ dataset (U.S., 2000-2022), we systematically compare statistical baselines, a fully supervised LSTM, and the Sundial foundation model under multiple evaluation protocols. We find that Sundial, in the zero-shot setting, can outperform a fully trained LSTM provided that the input context window is sufficiently long—specifically, when covering more than one or two full seasonal cycles. We show that a general-purpose foundation model can surpass specialized supervised models on remote-sensing time series prediction without any task-specific tuning. These results highlight the strong potential of pretrained time series foundation models to serve as effective plug-and-play forecasters in agricultural and environmental applications.
@@ Introduction
Leaf Area Index (LAI) is a critical biophysical variable widely used in agricultural and ecological monitoring, quantifying canopy structure and influencing crop growth. Beyond its central role in agronomic decision-making and yield estimation, accurate LAI forecasting is vital for understanding land-surface feedbacks, as vegetation significantly influences regional climate through biogeophysical processes (e.g., albedo and roughness) and biogeochemical cycles (Alessandri et al. 2017). However, practical forecasting is challenged by diverse cropping systems, varying environmental conditions, and limited labeled data. This motivates models that can generalize across locations and years. Furthermore, supervised models trained on heterogeneous datasets tend to learn ”global average” patterns and may miss site-specific phenological anomalies without extensive local fine-tuning.
Recent advances in time series foundation models, such as Sundial (Liu et al. 2025), offer a promising avenue for improving LAI forecasting. Unlike task-specific models, these large, pretrained networks leverage in-context learning, treating historical time series as ”prompts” to deduce underlying dynamics. This paradigm has the potential to adapt to new tasks and domains more efficiently than traditional supervised approaches.
@@
  • We analyze the context window threshold where zero-shot performance overtakes the supervised baseline, and discuss the distinct mechanisms at play: instance-level adaptation in foundation models versus dataset-level generalization in LSTMs.
@@
LAI Time Series Forecasting
Leaf Area Index (LAI) is a key indicator for monitoring vegetation health, phenology, and long-term ecosystem dynamics. Most LAI forecasting approaches treat the signal as a univariate time series characterized by strong seasonality, low-frequency variation, and limited annotated samples (Jiang et al. 2010). Classical methods include harmonic models, autoregressive formulations, and exponential-smoothing variants, which capture seasonal cycles but struggle with nonlinear responses to climate and land-cover variability (Winters 1960). Recent studies explore data-driven forecasting of vegetation indices, yet robust LAI prediction remains challenging due to heterogeneous temporal patterns and spatial variability across regions (Liu et al. 2022b).
@@
Classical LAI forecasting has long relied on statistical time series models such as autoregressive (AR) and exponential-smoothing formulations, which are simple, interpretable, and effective for short-term dynamics. However, these models are limited in their ability to capture nonlinear seasonal patterns and interannual variability.
Supervised machine-learning baselines—including random forests (Breiman 2001) and gradient boosting (Friedman 2001; Ke et al. 2017)—can incorporate richer auxiliary features when available, but offer only marginal improvements when restricted to univariate LAI histories. Recurrent neural networks, particularly LSTMs, remain strong supervised baselines for vegetation-index forecasting and are widely used in agronomic and ecological applications (Hochreiter and Schmidhuber 1997; Liu et al. 2022a). Despite their capacity to model nonlinear temporal dependencies, these methods require substantial labeled data and typically exhibit limited gains when training data is scarce or spatially heterogeneous.
Importantly, none of these approaches leverage cross-domain temporal knowledge. As a result, their generalization under long sequences, sparse supervision, or unseen spatial regions remains constrained—highlighting the value of pretrained time series foundation models, which can transfer temporal priors learned from large and diverse corpora.
Time Series Foundation Models
Large pretrained time series models have recently demonstrated strong zero-shot and few-shot transfer across diverse domains. Models such as TimesFM and Sundial learn universal temporal representations from mixed-modality corpora spanning climate, sensor networks, and economic indicators (Das et al. 2024; Liu et al. 2025; Ansari et al. 2024). These foundation models can leverage broad temporal priors—seasonality, trends, and multi-scale structure—reducing dependence on in-domain training data and enabling robust downstream forecasting.
@@
We conduct our experiments using the HiQ dataset (Yan et al. 2023), a large-scale benchmark for Leaf Area Index (LAI) prediction. HiQ reprocesses the satellite MODIS LAI product (Myneni et al. 2015) using a spatiotemporal information compositing algorithm, resulting in improved spatiotemporal consistency and reduced noise. The dataset provides high-quality global LAI time series spanning 2000–2022, with broad geographic coverage across diverse crop types and ecological regions.
In this work, we restrict our analysis to the continental United States (CONUS) subset. Although HiQ offers fully available global coverage, this choice is primarily driven by computational constraints rather than data availability or methodological limitations. We argue that the CONUS domain already encompasses a wide range of vegetation regimes and climatic conditions, making it a representative testbed for evaluating zero-shot LAI forecasting performance.
We use the HiQ product at 5-km spatial resolution, where each pixel contains an LAI time series sampled at an 8-day cadence, resulting in approximately 1,050 time steps per location.
@@ Methods
  • Trend: Extrapolates a fitted linear trend from the previous $k$ time points to predict future LAI values (Jensen et al. 2008).
  • ARIMA: For each test window, we fit an ARIMA(2,1,0)(2,1,0) model (Box and Pierce 1970; Seabold and Perktold 2010) to the observed LAI values to forecast the next value. ARIMA is chosen as a simple yet reasonable baseline for univariate time series; model fitting uses state-space estimation.
  • LSTM: A supervised Long Short-Term Memory model (Hochreiter and Schmidhuber 1997) trained on the entire training set to capture nonlinear dependencies in LAI time series. Our LSTM architecture consists of two stacked LSTM layers with 64 hidden units each, followed by a fully connected output layer. The model is trained using the Adam (Kingma 2014) optimizer with a learning rate of 0.001, batch size of 256, and early stopping based on validation loss.
@@
We formulate the LAI forecasting task as a conditional generation problem. Instead of fine-tuning model weights, we leverage the pre-trained capabilities of Sundial directly (Liu et al. 2025). The historical input window $X\in\mathbb{R}^{T_{\text{in}}}$ serves as a continuous context prompt, analogous to a text prompt in Large Language Models (LLMs). The model processes this prompt to deduce the underlying phenological dynamics (e.g., seasonality and trend) and generates the future sequence. This approach relies entirely on the model’s in-context learning ability acquired during pre-training, enabling a strictly zero-shot evaluation. We use the publicly available Sundial model checkpoint without any modifications or fine-tuning. For inference, we generate forecasts autoregressively for multi-step horizons.
@@
Metrics are aggregated over 100 randomly sampled locations to ensure computational efficiency. With approximately 100 rolling test windows per pixel, each reported metric represents an average across $\sim 10{,}000\times H$ prediction-ground truth pairs, where $H$ is the forecast horizon.
@@
Missing ratio RMSE increase over baseline (%)
1024 512 256 128 64
5% 8.2% 5.1% 1.5% 7.0% 0.8%
10% 10.2% 10.9% 2.1% 8.9% 4.2%
20% 25.6% 21.5% 12.8% 22.1% 7.5%
40% 57.1% 53.7% 42.3% 54.9% 36.7%
Table 2: Sundial robustness under random missing observations at forecast horizon $H=1$. Entries are relative RMSE increases (%) compared to the no-missing-data baseline.
@@
Robustness to Missing Data
To assess the robustness of Sundial to missing observations, we randomly mask a fraction of values in the input context at rates of 5%, 10%, and 20% without interpolation. Notably, Sundial is equipped with a native pipeline for handling NaN inputs. We report performance at forecast horizon $H=1$ as the relative RMSE increase (%) compared to the no-missing-data baseline (Table 2).
@@ Discussion
In this work, we conducted a systematic evaluation of the zero-shot capabilities of the Sundial time series foundation model for remote-sensing LAI forecasting. We found that, when provided with a sufficiently long historical context window, Sundial’s zero-shot performance can surpass that of a supervised LSTM model specifically trained for the task, challenging the conventional view that task-specific training guarantees superior performance. Our findings suggest that these general-purpose models already encode powerful temporal patterns and vegetation phenology, enabling them to serve as highly efficient and powerful plug-and-play forecasting tools without the need for any task-specific training.
@@
Acknowledgments
This work was partially supported by the U.S. National Science Foundation (NSF) under Grant AGS-2426655. We thank the anonymous reviewers for their constructive comments and suggestions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the NSF.